On a bright sunny day you (6ft) tall produce a shadow that is 10ft long. A tree in front of you produces a shadow 200ft long what is the hight of the tree?
Answer by stanbon(75887)

On a bright sunny day you (6ft) tall produce a shadow that is 10ft long. A tree in front of you produces a shadow 200ft long what is the height of the tree?
------------
x/200 = 6/10
x = 200(6/10)
x = 20*6
x = 120 ft (tree height)
=======================
Cheers,
Stan H.
=======================